Sharifi Hospital bases its budgets on patient-visits. The hospital's static budget for October appears below:

Budgeted number of patient-visits 9,500
Budgeted variable overhead costs:
Supplies (@$5.20 per patient-visit) $ 49,400
Laundry (@$8.20 per patient-visit) 77,900
Total variable overhead cost 127,300
Budgeted fixed overhead costs:
Wages and salaries 52,900
Occupancy costs 85,800
Total fixed overhead cost 138,700
Total budgeted overhead cost $ 266,000

The total overhead cost at an activity level of 10,200 patient-visits per month should be:

Multiple Choice

  • $276,220

  • $285,600

  • $266,000

  • $275,380

Teel Printing uses two measures of activity, press runs and book set-ups, in the cost formulas in its budgets and performance reports. The cost formula for wages and salaries is $4,350 per month plus $409 per press run plus $959 per book set-up. The company expected its activity in July to be 213 press runs and 120 book set-ups, but the actual activity was 210 press runs and 119 book set-ups. The actual cost for wages and salaries in July was $204,580.

The spending variance for wages and salaries in July would be closest to:

Multiple Choice

  • $219 F

  • $1,967 U

  • $219 U

  • $1,967 F

Rhed Kennel uses tenant-days as its measure of activity; an animal housed in the kennel for one day is counted as one tenant-day. During December, the kennel budgeted for 4,700 tenant-days, but its actual level of activity was 4,650 tenant-days. The kennel has provided the following data concerning the formulas used in its budgeting and its actual results for December:

Data used in budgeting:

Fixed element per month Variable element per tenant-day
Revenue - $ 35.90
Wages and salaries $ 4,200 $ 9.70
Food and supplies 2,400 14.90
Facility expenses 10,200 6.40
Administrative expenses 8,200 0.50
Total expenses $ 25,000 $ 31.50

Actual results for December:

Revenue $ 166,025
Wages and salaries $ 49,590
Food and supplies $ 70,555
Facility expenses $ 40,160
Administrative expenses $ 9,760

The activity variance for administrative expenses in December would be closest to:

Multiple Choice

  • $25 F

  • $25 U

  • $790 F

  • $790 U
Answer #1

1) Total overhead cost on 10200 visit = (127300/9500*10200)+138700 = 275380

So answer is d) $275380

2) Spending variance for wages and salaries = (4350+210*409+959*119)-204580 = 219 U

So answer is c) $219 U

3) Activity variance for administrative expense = (4700-4650)*.50 = 25 F

So answer is a) $25 F